# Unveiling quantum phase transitions from traps in variational quantum algorithms

### v1

@@ III Applications
The first term describes the interaction between neighbouring spins along the chain and the coupling constant $J$ quantifies the strength of this interaction. The second term represents the influence of a transverse magnetic field applied perpendicular to the direction of spin-spin interaction. The competition between these two terms leads to a quantum phase transition at the critical point $g=J$ [1, 6]. For $g<J$, the spin-spin interaction dominates, resulting in a phase where spins are ordered. For $g>J$, the transverse field dominates, leading to a disordered phase where spins align with the field. Usually, this transition is characterized by measuring the magnetization in the $x$-direction, $m_{x}=1/n\sum_{j=1}^{n}\sigma^{x}_{j}$, where $n$ is the number of spins in the system.
@@
Our approach employs the Hamiltonian variational ansatz, as detailed in Refs. [49, 50], parameterized by $\bm{\beta},\bm{\gamma}$:
@@
$U_{\text{TI-1D}}(\bm{\beta},\bm{\gamma})=\prod_{j=1}^{p}\exp\left(-i\frac{\beta_{j}}{2}\mathcal{H}_{x}\right)\exp\left(-i\frac{\gamma_{j}}{2}\mathcal{H}_{zz}\right),$ (16)
@@
where $\mathcal{H}_{zz}$ and $\mathcal{H}_{x}$ correspond to the sum of $\sigma^{z}_{j}\sigma^{z}_{j+1}$ and $\sigma^{x}_{j}$, respectively. As described in Sec. II.2, $\bm{\beta},\bm{\gamma}$ are represented by a Fourier series, with Fourier coefficients expressed as polynomials of $g$ up to the fourth order.
@@
The coupling strength $J$ is set to be 1. For the optimization process, the grid ranges from $g_{\text{min}}=0$ to $g_{\text{max}}=2$, with a sampling resolution of $\delta g=0.1$. The detection grid spans from $\hat{g}_{\text{min}}=0.57$ to $\hat{g}_{\text{max}}=1.13$, with a finer sampling resolution of $\delta\hat{g}=0.001$. As long as $n\geq 2p+2$, the optimized local observable expectation values are ensured to remain consistent for infinitely large $n$. Beginning with $p=1$ and $n=4$, we sample initial values for $\bm{\zeta}$ and $\bm{\eta}$ from a narrow range $(-10^{-3},10^{-3})$, then we iteratively adjust their values to minimize the energy sum over $\mathcal{G}_{\text{opt}}$. We found that different samples of the initial values of $\bm{\zeta}$ and $\bm{\eta}$ within this near-zero range do not lead to significant differences in the outcomes. Hence, we present results based on a singular random sample in subsequent analyses.
@@
$\left.\sigma_{i}^{x}\sigma_{j}^{y},\sigma_{i}^{y}\sigma_{j}^{z},\sigma_{i}^{z}\sigma_{j}^{x}\right\}_{i\neq j}.$
@@
thereby extending $\tilde{\bm{\beta}}$ and $\tilde{\bm{\gamma}}$ with zero initial values to accommodate the expanded circuit and system size. This procedure is iteratively executed in numerical simulations, enabling the prediction of critical points $\tilde{g}_{c}$ for various values of $p$. The recursive optimization strategy is depicted in Fig. 4. Instead of setting $n=2p$ for producing the exact ground states [50], our simulations focuses on configurations where $n\geq 2p+2$—practically, $n=2p+2$—aiming to understand the finite-depth effects and explore the applications of finite-depth extrapolation discussed in Section II.3.
@@
$\mathcal{F}(g)=|\langle\psi(\bm{\zeta},\bm{\eta};g)|\psi_{\mathrm{gs}}\rangle|^{2},$ (19)
for $p=2,3,\cdots,7$ and $n=18$ across the detection grid $\mathcal{G}_{\text{det}}$. Fidelities at the critical point $g_{c}=1$ are recorded as 0.58, 0.68, 0.75, 0.80, 0.84, and 0.87, respectively. Near $g\approx 0$, fidelities exhibit considerable volatility, a phenomenon attributed to the minimal spectral gap in the region where $g$ is close but not equal to zero. Despite the optimized quantum states achieving relatively stable low energy, the fidelity to the ground state remains inconsistent. Nevertheless this region is distant from $g_{c}=1$ and minimally impacts our findings. For $g\gg 1$, even at $p=2$, the fidelities significantly increase, as the ground states are close to the initial product state. The inset of panel (a) displays the derivative of fidelity, $\partial\mathcal{F}/\partial g$, as a function of $g$, revealing that the peak does not align with the thermodynamic critical point and shifts away from it as $p$ increases. Panel (b) presents the LASSO regression’s loss landscape, highlighting the minimum point which correlates with the anticipated phase transition critical point $\tilde{g}_{c}$. Utilizing a relatively large regularization parameter $\lambda$ sharpens the valleys within the loss landscape, leading to the selection of a singular observable as the order parameter at the valley’s lowest point. The LASSO selected order parameter here is
@@
The exponential decay in the error between the estimated and the exact critical point as a function of the circuit depth $p$, as predicted by our algorithm and by $m_{x}$, is shown in Fig. 5(c,d). This further supports the application of exponential extrapolation techniques, discussed in Section II.3, to accurately determine the critical point in the thermodynamic limit. By applying Eq. (13) to the finite-$p$ critical points obtained from $m_{x}$, we obtain an extrapolated critical point $\tilde{g}_{c}(p\rightarrow\infty)=0.921$, with a mean square fitting error of 0.004. Moreover, using the finite-depth extrapolation technique in combination with our algorithm predicts $\tilde{g}_{c}(p\rightarrow\infty)=0.995$, yielding a mean square fitting error of 0.002, closely approximating the theoretical value of $g_{c}=1$. Compared with prior works [20, 24], our approach achieves stable and precise estimations of the critical point even with notably shallow circuits. For further evaluation, we applied polynomial fitting using the equation $\tilde{g}_{c}(p)=g_{c}(p\rightarrow\infty)+cp^{-\nu}$. The extrapolated $g_{c}(p\rightarrow\infty)$ values for $m_{x}$ and our algorithm are $\tilde{g}_{c}(p\rightarrow\infty)=1.566\pm 0.003$ and $\tilde{g}_{c}(p\rightarrow\infty)=2.545\pm 0.004$, respectively, both significantly deviating from the theoretical value of $g_{c}=1$, further corroborating Observation 1. Furthermore, we compute the Pearson correlation coefficient $r$ between the estimation error $|\tilde{g}_{c}-g_{c}|$ and the fidelity at the critical point $\mathcal{F}(g_{c})$, defined as [62]:
@@
III.2 2D Transverse-field Ising Model
Now we analyse a 2D TFIM consisting of a $3\times 3$ qubit grid with open boundary conditions, described by the Hamiltonian
@@
$U_{\text{TI-2D}}(\bm{\beta},\bm{\gamma})=\prod_{j=1}^{p}$ $\exp\left(-i\frac{\beta_{j}}{2}\mathcal{H}_{x}\right)\exp\left(-i\frac{\beta^{\prime}_{j}}{2}\mathcal{H}^{\prime}_{x}\right)$ (24)
@@
To facilitate the estimation of expectation values for all Pauli terms in $\mathcal{P}_{f}$, which consist of weight one and two operators up to rotation and reflection symmetries, we implement four distinct measurement protocols, detailed in Fig. 6(b). For qubits labeled with $X$, $Y$, and $Z$, pre-measurement single-qubit unitaries are applied to rotate the measurement basis accordingly, followed by projective measurements. This methodology ensures comprehensive analysis and measurement across the qubit array.
Before implementing our algorithm across a range of $g$ values, we examine two metrics which have been used to analyse quantum phase transitions in the literature [63, 64, 65, 66]: the relative spectral gap, defined as $(E_{1}-E_{\text{gs}})/E_{\text{gs}}$ where $E_{\text{gs}}$ represents the ground state energy and $E_{1}$ denotes the first excited state energy of $\mathcal{H}_{\text{TI-2D}}$, and the ground state entanglement entropy, with a focus on the entanglement between the central qubit, $Q_{4}$, and its complement, $Q^{c}_{4}$, within the system. This is expressed as
@@
In our exploration, the relative spectral gap and the entanglement entropy serve as indicators of shifts in the global properties of the ground state, as opposed to merely reflecting changes in local order parameters. Consequently, in comparison to a local order parameter such as the magnetization, these metrics are less affected by finite-size effects and emerge as more sensitive probes for detecting phase transitions within small systems. By computing the derivatives of these metrics, we identify peaks at $1.84$ and $1.68$, respectively, as illustrated in Fig. 6(c). These peaks hint at a likely phase transition with critical point located near $1.68\sim 1.84$. In contrast, the derivative of the $x$-magnetization, $m_{x}=1/n\sum_{j=1}^{n}\sigma^{x}_{j}$, exhibits a sharp decline at $1.34$, markedly deviating from the aforementioned range. Numerical analyses have determined the phase transition point for the 2D TFIM in the thermodynamic limit to be approximately $3.04438$ [67, 68]. In the infinite-size limit, predictions of the critical point from relative energy gap, entanglement entropy, and $x$-magnetization are expected to converge to this value. However, due to hardware constraints, our analysis is limited to small system sizes only. Here, estimates of the critical point from the peaks in the relative energy gap and entanglement entropy are closer to the exact value and confirm the enhanced sensitivity of these probes with respect to a local order parameter. Since obtaining a precise estimate of the critical point would require a finite-size (or depth) extrapolation, which is beyond the scope of this work, in the following we will assume the range $1.68\sim 1.84$ as a benchmark for our protocol.
@@
Both numerical and experimental investigations reveal a loss minimum at 1.6, as depicted in Fig. 6(d), offering a refinement over estimates derived from $m_{x}$. This also show that our algorithm can accurately extract phase transition information without necessitating error mitigation. By integrating symmetry considerations into our assessment, we find the following learned order parameter
@@
The inset of Fig. 6(d) showcases the derivative of $\langle\mathcal{O}_{\text{ml}}\rangle$ with respect to $g$, calculated using next-nearest values. Additionally, we assess the impact of hardware noise on $\mathcal{P}_{f}$, analyzed under the quasi-GD noise approximation outlined in Theorem 1. For the purposes of our algorithm, the influence of hardware noise on $\mathcal{P}_{f}$ resembles a global depolarizing noise channel $\Lambda_{\epsilon}(\cdot)$ with a noise rate of $\epsilon=0.773$. The average discrepancy between the ideal and the linearly rescaled noisy features is $0.047$.
@@
resulting in a significantly less accurate estimate. The inset in Fig. 6(e) shows the relationship between $\tilde{g}_{c}$ and the noise rate for both amplitude damping and depolarizing channels, with crossover points at $\epsilon^{(\text{AD})}=0.064$ and $\epsilon^{(\text{DP})}=0.028$, respectively.
III.3 Extended Su–Schrieffer–Heeger Model

### v2

@@ II Results
The first term describes the interaction between neighbouring spins along the chain and the coupling constant $J$ quantifies the strength of this interaction. The second term represents the influence of a transverse magnetic field applied perpendicular to the direction of spin-spin interaction. The competition between these two terms leads to a quantum phase transition at the critical point $g=J$ [1, 6]. For $g<J$, the spin-spin interaction dominates, resulting in a phase where spins are ordered. For $g>J$, the transverse field dominates, leading to a disordered phase where spins align with the field. Usually, this transition is characterized by measuring the magnetization in the $x$-direction, $m_{x}=1/n\sum_{j=1}^{n}\sigma^{x}_{j}$, where $n$ is the number of spins in the system. In what follows, the coupling strength $J$ is set to be 1.
@@
Our approach employs the Hamiltonian variational ansatz, as detailed in Refs. [63, 64], parameterized by $\boldsymbol{\beta},\boldsymbol{\gamma}$ (see Section IV.2):
@@
$U_{\text{TI-1D}}(\boldsymbol{\beta},\boldsymbol{\gamma})=\prod_{j=1}^{p}\exp\left(-i\frac{\beta_{j}}{2}\mathcal{H}_{x}\right)\exp\left(-i\frac{\gamma_{j}}{2}\mathcal{H}_{zz}\right),$ (2)
@@
where $\mathcal{H}_{zz}$ and $\mathcal{H}_{x}$ correspond to the sum of $\sigma^{z}_{j}\sigma^{z}_{j+1}$ and $\sigma^{x}_{j}$, respectively. As described in section IV.2, $\boldsymbol{\beta},\boldsymbol{\gamma}$ are represented by a Fourier series, with Fourier coefficients $\tilde{\boldsymbol{\beta}}$ and $\tilde{\boldsymbol{\gamma}}$ expressed as polynomials of $g$ up to the fourth order,
$\tilde{\gamma}_{k}=\sum_{j=0}^{4}\zeta_{j,k}g^{j},\quad\tilde{\beta}_{k}=\sum_{j=0}^{4}\eta_{j,k}g^{j}.$ (3)
The vectors $\boldsymbol{\zeta}$ and $\boldsymbol{\eta}$ are then optimized to minimize the energy function sum
$\mathcal{E}_{\text{opt}}=\sum_{g\in\mathcal{G}_{\text{opt}}}\operatorname{Tr}\left(\mathcal{H}_{\text{TI-1D}}(g)\rho(\boldsymbol{\zeta},\boldsymbol{\eta};g)\right),$ (4)
where $\rho(\boldsymbol{\zeta},\boldsymbol{\eta};g)$ are the optimized states and $\mathcal{G}_{\mathrm{opt}}=\{g_{\mathrm{min}},g_{\mathrm{min}}+\delta g,...,g_{\mathrm{max}}\}$, with $\delta g$ the optimization sampling resolution, is the optimization grid. The global optimization approach we adopt here ensures better stability and allows us to generate optimized states for any value of $g$. This is particularly useful in training the machine learning algorithm for the detection the system’s quantum phase transitions, where we prepare optimized states $\rho(\boldsymbol{\zeta},\boldsymbol{\eta};g)$ for all the values of $g$ in the detection grid $\mathcal{G}_{\mathrm{det}}=\{\hat{g}_{\mathrm{min}},\hat{g}_{\mathrm{min}}+\delta\hat{g},...,\hat{g}_{\mathrm{max}}\}$, with $\delta\hat{g}$ the detection sampling resolution. See section IV.2 for more details.
@@
For the optimization process of the 1D TFIM, we choose $g_{\text{min}}=0$ to $g_{\text{max}}=2$, with a sampling resolution of $\delta g=0.1$. The detection grid spans from $\hat{g}_{\text{min}}=0.57$ to $\hat{g}_{\text{max}}=1.13$, with a finer sampling resolution of $\delta\hat{g}=0.001$. As long as $n\geq 2p+2$, the optimized local observable expectation values are ensured to remain consistent for infinitely large $n$. Beginning with $p=1$ and $n=4$, we sample initial values for $\boldsymbol{\zeta}$ and $\boldsymbol{\eta}$ from a narrow range $(-10^{-3},10^{-3})$, then we iteratively adjust their values to minimize the energy sum over $\mathcal{G}_{\text{opt}}$. We found that different samples of the initial values of $\boldsymbol{\zeta}$ and $\boldsymbol{\eta}$ within this near-zero range do not lead to significant differences in the outcomes. Hence, we present results based on a singular random sample in subsequent analyses.
@@
$\sigma_{i}^{x}\sigma_{j}^{y},\sigma_{i}^{y}\sigma_{j}^{z},\sigma_{i}^{z}\sigma_{j}^{x}\big{\}}_{i\neq j}.$
@@
thereby extending $\tilde{\boldsymbol{\beta}}$ and $\tilde{\boldsymbol{\gamma}}$ with zero initial values to accommodate the expanded circuit and system size. This procedure is iteratively executed in numerical simulations, enabling the prediction of critical points $\tilde{g}_{c}$ for various values of $p$. The recursive optimization strategy is depicted in Fig. 2.
Instead of setting $n=2p$ for producing the exact ground states [64], our simulations focuses on configurations where $n\geq 2p+2$—practically, $n=2p+2$—aiming to understand the finite-depth effects and explore the applications of finite-depth extrapolation discussed in Section IV.3.
@@
$\mathcal{F}(g)=|\langle\psi(\boldsymbol{\zeta},\boldsymbol{\eta};g)|\psi_{\mathrm{gs}}\rangle|^{2},$ (7)
for $p=2,3,\cdots,7$ and $n=18$ across the detection grid $\mathcal{G}_{\text{det}}$. Fidelities at the critical point $g_{c}=1$ are recorded as 0.58, 0.68, 0.75, 0.80, 0.84, and 0.87, respectively. Near $g\approx 0$, fidelities fluctuate considerably due to the small spectral gap present in this region, where optimization struggles to consistently identify states close to the true ground state. Despite the optimized quantum states achieving relatively stable low energy, the fidelity to the ground state remains inconsistent. Nevertheless this region is distant from $g_{c}=1$ and minimally impacts our findings. For $g\gg 1$, even at $p=2$, the fidelities significantly increase, as the ground states are close to the initial product state. The inset of panel (a) displays the derivative of fidelity, $\partial\mathcal{F}/\partial g$, as a function of $g$, revealing that the peak does not align with the thermodynamic critical point and shifts away from it as $p$ increases. Panel (b) presents the LASSO regression’s loss landscape, highlighting the minimum point which correlates with the anticipated phase transition critical point $\tilde{g}_{c}$. Utilizing a relatively large regularization parameter $\lambda$ sharpens the valleys within the loss landscape, leading to the selection of a singular observable as the order parameter at the valley’s lowest point. The LASSO selected order parameter here is
@@
The exponential decay in the error between the estimated and the exact critical point as a function of the circuit depth $p$, as predicted by our algorithm and by $m_{x}$, is shown in Fig. 3(c,d). This suggests that, under certain assumptions (see Observation 1 in section IV.3), the location of the critical point in the thermodynamic limit, $g_{c}$, can be predicted by applying an exponential extrapolation approach to the results for finite depth $p$, $\tilde{g}_{c}(p)$, i.e., $\tilde{g}_{c}(p)=g_{c}(p\rightarrow\infty)+ce^{-\nu p}$. By applying the equation above to the finite-$p$ critical points obtained from $m_{x}$, we obtain an extrapolated critical point $\tilde{g}_{c}(p\rightarrow\infty)=0.921$, with a mean square fitting error of 0.004. Moreover, using the finite-depth extrapolation technique in combination with our algorithm predicts $\tilde{g}_{c}(p\rightarrow\infty)=0.995$, yielding a mean square fitting error of 0.002, closely approximating the theoretical value of $g_{c}=1$. Compared with prior works [21, 25], our approach achieves stable and precise estimations of the critical point even with notably shallow circuits. For further evaluation, we applied polynomial fitting using the equation $\tilde{g}_{c}(p)=g_{c}(p\rightarrow\infty)+cp^{-\nu}$. The extrapolated $g_{c}(p\rightarrow\infty)$ values for $m_{x}$ and our algorithm are $\tilde{g}_{c}(p\rightarrow\infty)=1.566\pm 0.003$ and $\tilde{g}_{c}(p\rightarrow\infty)=2.545\pm 0.004$, respectively, both significantly deviating from the theoretical value of $g_{c}=1$, further corroborating the exponential convergence of $\tilde{g}_{c}(p)$ to the thermodynamic limit value. Furthermore, we compute the Pearson correlation coefficient $r$ between the estimation error $|\tilde{g}_{c}-g_{c}|$ and the fidelity at the critical point $\mathcal{F}(g_{c})$, defined as [66]:
@@
Figure 4: Estimation of critical points using various learned order parameters in the 1D transverse-field Ising model. (a) Results for optimized states at circuit depth $p=4$, with the average of $\mathcal{O}_{\mathrm{xx}}(d)$ estimating the critical point as $\tilde{g}_{c}=0.83\pm 0.05$, the average of $\mathcal{O}_{\mathrm{zz}}(d)$ as $\tilde{g}_{c}=0.96\pm 0.03$, and the average of $\mathcal{O}_{\mathrm{yz}}(d)$ as $\tilde{g}_{c}=0.97\pm 0.06$. (b) Results for circuit depth $p=7$. The average of $\mathcal{O}_{\mathrm{xx}}(d)$ estimates the critical point as $\tilde{g}_{c}=0.94\pm 0.03$, the average of $\mathcal{O}_{\mathrm{zz}}(d)$ as $\tilde{g}_{c}=1.00\pm 0.02$, and the average of $\mathcal{O}_{\mathrm{yz}}(d)$ as $\tilde{g}_{c}=0.99\pm 0.04$. In both panels, dashed lines indicate the average critical point estimates across different $d$ values while the dotted lines display the critical point estimate obtained from the magnetization $m_{x}$.
@@
As we will discuss in section IV.2, our algorithm can identify the most dominant feature that serves as an order parameter to characterize the phase transition. As an interesting extension, we can iteratively refine this process by manually removing the currently identified feature, re-running the algorithm to find the next dominant feature, and repeating this process until no stable central-minima loss valley is observed. This approach allows us to identify a series of order parameters that provide estimates of the critical point. We implement this iterative approach for optimized states from quantum circuits with depths $p=4$ and $p=7$, identifying the following symmetric order parameters:
$\mathcal{O}_{\mathrm{xx}}(d)=\frac{1}{n}\sum_{j}\sigma^{x}_{j}\sigma^{x}_{j+d},$ (11)
$\mathcal{O}_{\mathrm{zz}}(d)=\frac{1}{n}\sum_{j}\sigma^{z}_{j}\sigma^{z}_{j+d}$
where $d=1,2,\ldots,p+1$, and
$\mathcal{O}_{\mathrm{yz}}(d)=\frac{1}{2n}\sum_{j}(\sigma^{y}_{i}\sigma^{z}_{j+d}+\sigma^{z}_{j}\sigma^{y}_{j+d})$ (12)
where $d=2,3,\ldots,p+1$. The estimated critical points obtained from these learned order parameters over all the possible values of $d$ are shown in Fig. 4. Each order parameter provides a more precise estimate of the critical point compared to traditional $x$-magnetization. Notably, even for these states with fidelities of only 0.75 ($p=4$) and 0.87 ($p=7$) at the critical point, $\mathcal{O}_{\mathrm{zz}}$ and $\mathcal{O}_{\mathrm{yz}}$ provides an accurate estimate of the critical point when $d\geq 4$: $0.98\lesssim\tilde{g}_{c}\lesssim 1.03$ without the need for extrapolation. Additionally, we notice that the difference between the estimated critical point $\tilde{g}_{c}$ and the theoretical value $g_{c}=1$ tends to diminish as $d$ increases. Furthermore, for relatively small values of $d$, extrapolation in $p$ can be employed to refine the critical point estimation.
Additionally, we investigate the influence of noise on the optimization of the variational parameters. Our numerical simulations on the 1D TFIM (detailed in the Supplementary Material (SM)) reveal that even when the optimization is performed on noisy quantum circuits—where each 2-qubit gate is immediately followed by a local depolarizing noise channel with error rate $\epsilon^{(\text{DP})}=0.01$—the resulting optimized parameters are only marginally affected. Both the output fidelity as a function of $g$ and the LASSO loss landscapes for circuit depths $p=2$ and $p=3$ show only slight degradations compared to the noiseless case. Notably, the learned order parameter remains $\mathcal{O}_{\mathrm{ml}}=(1/n)\sum_{j}\sigma_{j}^{x}\sigma_{j+2}^{x}$, and the critical point estimates derived from the loss landscapes differ only minimally. These results indicate that our algorithm is robust to moderate hardware noise during the optimization process.
@@
II.2 2D Transverse-field Ising Model
Now we first analyse a 2D TFIM consisting of a $3\times 3$ qubit grid with open boundary conditions, described by the Hamiltonian
@@
$U_{\text{TI-2D}}(\boldsymbol{\beta},\boldsymbol{\gamma})=\prod_{j=1}^{p}$ $\exp\left(-i\frac{\beta_{j}}{2}\mathcal{H}_{x}\right)\exp\left(-i\frac{\beta^{\prime}_{j}}{2}\mathcal{H}^{\prime}_{x}\right)$ (14)
@@
To facilitate the estimation of expectation values for most Pauli terms in $\mathcal{P}_{f}$, which consist of weight one and two operators up to rotation and reflection symmetries, we implement four distinct measurement protocols, detailed in Fig. 5(b). For qubits labeled with $X$, $Y$, and $Z$, pre-measurement single-qubit unitaries are applied to rotate the measurement basis accordingly, followed by projective measurements. This methodology ensures comprehensive analysis and measurement across the qubit array.
Before implementing our algorithm across a range of $g$ values, we examine two metrics which have been used to analyse quantum phase transitions in the literature [67, 68, 69, 70]: the relative spectral gap, defined as $|(E_{1}-E_{\text{gs}})/E_{\text{gs}}|$ where $E_{\text{gs}}$ represents the ground state energy and $E_{1}$ denotes the first excited state energy of $\mathcal{H}_{\text{TI-2D}}$, and the ground state entanglement entropy, with a focus on the entanglement between the central qubit, $Q_{4}$, and its complement, $Q^{c}_{4}$, within the system. This is expressed as
@@
In our exploration, the relative spectral gap and the entanglement entropy serve as indicators of shifts in the global properties of the ground state, as opposed to merely reflecting changes in local order parameters. Consequently, in comparison to a local order parameter such as the magnetization, these metrics are less affected by finite-size effects and emerge as more sensitive probes for detecting phase transitions within small systems. By computing the derivatives of these metrics, we identify peaks at $1.84$ and $1.68$, respectively, as illustrated in Fig. 5(c). These peaks hint at a likely phase transition with critical point located near $1.68\sim 1.84$. In contrast, the derivative of the $x$-magnetization, $m_{x}=1/n\sum_{j=1}^{n}\sigma^{x}_{j}$, exhibits a sharp decline at $1.34$, markedly deviating from the aforementioned range. Previous numerical studies have determined that the 2D TFIM undergoes a quantum phase transition at a critical point of approximately $3.04438$ in the thermodynamic limit [71, 72]. In the infinite-size limit, predictions of the critical point from relative energy gap, entanglement entropy, and $x$-magnetization are expected to converge to this value. However, owing to current hardware constraints, our analysis in this part is confined to a $3\times 3$ qubit system. To further validate our protocol, we extend our numerical investigations to larger system sizes with periodic boundary conditions in the SM. For the $3\times 3$ system considered here, estimates of the critical point from the peaks in the relative energy gap and entanglement entropy are also closer to the exact value and confirm the enhanced sensitivity of these probes with respect to a local order parameter. Since obtaining a precise estimate of the critical point would require a finite-size (or depth) extrapolation, which is beyond the scope of this work, in the following we will assume the range $1.68\sim 1.84$ as a benchmark for our protocol. If an order parameter yields an estimate that is closer to, or even surpasses, this benchmark—indicating that it converges more rapidly and is less affected by finite-size effects—then we consider that order parameter to be superior.
@@
Both numerical and experimental investigations reveal a loss minimum at 1.6, as depicted in Fig. 5(d), offering a refinement over estimates derived from $m_{x}$. This also shows that our algorithm can accurately extract phase transition information without necessitating error mitigation. By integrating symmetry considerations into our assessment, we find the following learned order parameter
@@
The inset of Fig. 5(d) showcases the derivative of $\langle\mathcal{O}_{\text{ml}}\rangle$ with respect to $g$, calculated using next-nearest values. Additionally, we assess the impact of hardware noise on $\mathcal{P}_{f}$, analysed under the quasi-GD noise approximation outlined in Theorem 1. For the purposes of our algorithm, the influence of hardware noise on $\mathcal{P}_{f}$ resembles a global depolarizing noise channel $\Lambda_{\epsilon}(\cdot)$ with a noise rate of $\epsilon=0.773$. The average discrepancy between the ideal and the linearly rescaled noisy features is $0.047$.
@@
resulting in a significantly less accurate estimate. The inset in Fig. 5(e) shows the relationship between $\tilde{g}_{c}$ and the noise rate for both amplitude damping and depolarizing channels, with crossover points at $\epsilon^{(\text{AD})}=0.064$ and $\epsilon^{(\text{DP})}=0.028$, respectively. In summary, our results demonstrate that the algorithm exhibits robust performance for small to moderate noise levels—i.e., the predicted phase transition point remains nearly unchanged relative to the ideal case. The observed abrupt shift at higher noise rates is consistent with the expectation that substantial noise will fundamentally alter the quantum state, thereby affecting the performance of any algorithm.
As discussed in the previous section, the iterative removal of identified dominant order parameters from the feature vectors facilitates the discovery of multiple additional order parameters. Table 2 presents the most dominant order parameter and some additional order parameters alongside the corresponding simulation and experimental results, showing a generally consistent alignment between the two.
Additional Learned Order Parameters Sim. Exp.
$(1/4)(\sigma^{x}_{1}\sigma^{x}_{4}+\sigma^{x}_{3}\sigma^{x}_{4}+\sigma^{x}_{4}\sigma^{x}_{5}+\sigma^{x}_{4}\sigma^{x}_{7})$ 1.60 1.6
$(1/2)(\sigma^{z}_{0}\sigma^{z}_{8}+\sigma^{z}_{2}\sigma^{z}_{6})$ 1.29 1.3
$(1/2)(\sigma^{z}_{1}\sigma^{z}_{7}+\sigma^{z}_{3}\sigma^{z}_{5})$ 1.47 1.3
$(1/4)(\sigma^{z}_{1}\sigma^{z}_{3}+\sigma^{z}_{1}\sigma^{z}_{5}+\sigma^{z}_{3}\sigma^{z}_{7}+\sigma^{z}_{5}\sigma^{z}_{7})$ 1.27 1.4
$(1/4)(\sigma^{z}_{0}\sigma^{z}_{2}+\sigma^{z}_{0}\sigma^{z}_{6}+\sigma^{z}_{2}\sigma^{z}_{8}+\sigma^{z}_{6}\sigma^{z}_{8})$ 1.27 1.3
$(1/2)(\sigma^{x}_{1}\sigma^{x}_{7}+\sigma^{x}_{3}\sigma^{x}_{5})$ 1.52 1.8
Table 2: Comparison of predicted critical points from simulation and experimental results using different learned order parameters.
In the SM, we investigate the critical boundary of 2D TFIM and the performance of our protocol for larger qubit grids. Applying our algorithm to the $4\times 4$ system yields an estimated critical point of approximately 2.45, which is significantly closer to the thermodynamic limit value.
II.3 Extended Su–Schrieffer–Heeger Model